OpenStudy (anonymous):

7x+11y=5462, solve for x and y

OpenStudy (anonymous):

you cant

OpenStudy (anonymous):

i got the answer: x=780.29-1.57y y=496.55-0.64x

OpenStudy (anonymous):

the answer does not seem right to me

OpenStudy (anonymous):

you simply rewrote the equation in a different way. It doesn't mean you solved for x and y